OpenStudy (anonymous):

18. For the function rule y = -2x - 3, provide the x-y coordinates of the y-intercept and the x-intercept

OpenStudy (anonymous):

y intercept is when x = 0 So plug in for x in y= -2x - 3 You get y = -2(0) - 3 = -3 So when x = 0, y = -3 that is (0,-3) Similarly, x intercept is when y = 0 So again, plug in 0 for y to get 0 = -2x - 3 so that 2x = -3 and x = -3/2 Thus, you have (-3/2 , 0) as the second coordinate

OpenStudy (anonymous):

is that the answer . sorry im just a little confused.

OpenStudy (anonymous):

(0,-3) and (-3/2,0)

OpenStudy (anonymous):

thank u soo much :D

OpenStudy (anonymous):

No problem!